# On the Fundamental Impossibility of Hallucination Control in Large Language Models

### v1

@@ 6 The Impossibility Theorem
We now present and prove our main result demonstrating the fundamental impossibility of perfect hallucination control.
@@
By definition of a non-trivial query $q$, given a random knowledge profile, such that
$P(\theta_{1},\dots,\theta_{n})=\prod_{i=1}^{n}P_{i}(\theta_{i}),$
the hallucination cost is minimized only when $\theta=(\theta_{1},\dots,\theta_{n})$ is integrated into response $r^{*}=g(s^{*}(\theta))$. If every agent follows its truthful strategy $s^{*}(\theta)$, the union of the knowledge elements contained in the agents’ inference paths reconstructs the set of all knowledge that is computationally accessible, i.e.,
$\Psi(\theta)\;=\;\bigcup_{i=1}^{n}K\!\bigl{(}s_{i}^{*}(\theta_{i})\bigr{)}.$ (23)
That means that for any proper subset of agents $A\subset\mathcal{A}$, relevant contribution reduces hallucination cost:
$H(g((\mathbf{0}_{-A},s^{*}_{A})),q)>H(g(s^{*}(\theta)),q),$ (24)
where $s^{*}_{A}$ represents truthful strategies from agents in subset $A$ and $\mathbf{0}_{-A}$ represents null contributions from other agents.
Since we have assumed independence of private-knowledge, by the Green-Laffont theorem every truthful response mechanism $\mathcal{M}$ must use payments of the form:
$p_{i}(s,\theta)=p_{i}(s)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j}),$ (25)
where $h_{i}(s_{-i})$ is arbitrary functions that do not depend on agent $i$’s strategy. This is a fundamental requirement for ensuring truthfulness.
To see that the theorem applies to our LLM inference auction model, suppose there exists another payment function $\bar{p}_{i}$ that leads to truthful responses and let us define:
$\Delta p_{i}(s)=\bar{p}_{i}(s)-\Big{(}h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})\Big{)}.$
Then, for truthful $s_{i}=s_{i}(\theta_{i})$ and any $\bar{s}_{i}=\bar{s}_{i}(\theta)$:
$v_{i}(g((s_{i},s_{-i})),\theta_{i})-\bar{p}_{i}((s_{i},s_{-i}))\geq v_{i}(g((\bar{s}_{i},s_{-i})),\theta_{i})-\bar{p}_{i}((\bar{s}_{i},s_{-i})).$
Substituting the definition of $\Delta p_{i}$ and simplifying, we get:
$\sum_{i=1}^{n}v_{i}(g((s_{i},s_{-i})),\theta_{i})-\Delta p_{i}((s_{i},s_{-i}))\geq\sum_{i=1}^{n}v_{i}(g((\bar{s}_{i},s_{-i})),\theta_{i})-\Delta p_{i}((\bar{s}_{i},s_{-i})).$
It follows from the knowledge-constrained optimality that $\Delta p_{i}((s_{i},s_{-i}))\leq\Delta p_{i}((\bar{s}_{i},s_{-i}))$ for all $s_{i},\bar{s}_{i}$. However, by the symmetry of argument, for truthful $\bar{s}_{i}=\bar{s}_{i}(\bar{\theta}_{i})$ and any $s_{i}=s_{i}(\bar{\theta}_{i})$ we also get $\Delta p_{i}((\bar{s}_{i},s_{-i}))\leq\Delta p_{i}((s_{i},s_{-i})).$ Therefore,
$\Delta p_{i}((s_{i},s_{-i}))=\Delta p_{i}((\bar{s}_{i},s_{-i}))$
for all $s_{i},\bar{s}_{i}$. Expanding and simplifying yields:
$\bar{p}_{i}((s_{i},s_{-i}))-\bar{p}_{i}((\bar{s}_{i},s_{-i}))=-\sum_{j\neq i}v_{j}(g((s_{i},s_{-i}),\theta_{j})-\Big{(}-\sum_{j\neq i}v_{j}(g((\bar{s}_{i},s_{-i})),\theta_{j})\Big{)},$
which proves $\bar{p}_{i}(s,\theta)$ must have the form of $p(s,\theta)$.
The payment structure above is necessary to controls the inference process and incentivizes the agents to minimize the hallucination cost, contributing to knowledge-constrained optimality of response. Since non-trivial queries require integrated knowledge from multiple agents, it follows that each contributing agent $i$ reduces that cost proportionally to:
$\Delta_{i}=\sum_{j\neq i}v_{j}(g((s_{-i},0_{i})),\theta_{j})-\sum_{j\neq i}v_{j}(g(s),\theta_{j}).$
Since we also require that all agents reveal relevant knowledge in a non-hallucinatory way with $s^{*}_{i}=s^{*}_{i}(\theta_{i})$ , we conclude that:
$\Delta_{i}<0\quad\text{whenever}\quad u_{i}(s^{*}(\theta),\theta_{i})=v_{i}(g(s^{*}(\theta)),\theta_{i})-p_{i}(s^{*},\theta)\geq 0.$
But that directly contradicts the information conservation requirement. Indeed, any payment structure that ensures non-hallucinatory truthfulness while optimizing responses for knowledge-constrained optimality will make each payment $p_{i}(s^{*},\theta)$ equal to $\Delta_{i}<0$ near $s^{*}=s^{*}(\theta)$, thereby violating the semantic information conservation principle:
$\sum_{i=1}^{n}p_{i}(s^{*},\theta)<0.$ (26)
As demonstrated earlier, the semantic information conservation equation $\sum_{i=1}^{n}p_{i}(s^{*},\theta)=0$ is necessary and sufficient for every truthful response to be of bounded response creativity. Since that equation is not satisfied when truthful responses use all relevant knowledge under any form of knowledge-constrained optimal inference, the responses must introduce some creativity (like e.g. lucky hallucinations).
To satisfy information conservation, we can try modifying the $h_{i}$ functions. Let us consider:
$\hat{p}_{i}(s,\theta)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})+c_{i}=\Delta_{i}+c_{i},$ (27)
where $c_{i}$ are constants chosen so that $\sum_{i=1}^{n}\hat{p}_{i}(s,\theta)=0$. For this to hold:
$\sum_{i=1}^{n}c_{i}=-\sum_{i=1}^{n}\Delta_{i}>0.$ (28)
This means that $c_{i}>0$ for some agents. For those with small contributions (which exist in LLMs supporting the superposition phenomenon) the modified payment would increase:
$\hat{p}_{i}(s^{*},\theta)=\Delta_{i}+c_{i}>p(s^{*},\theta).$ (29)
Therefore, it the setting considered, we can no longer guarantee relevant knowledge revelation. There may exist agent $i$ with small value $v_{i}$ and $\Delta_{i}$ for which:
$u_{i}(s^{*}(\theta),\theta_{i})=v_{i}(g(s^{*}(\theta)),\theta_{i})-\Delta_{i}-c_{i}<0,$ (30)
when $c_{i}>0$. Marginal relevant contributions will be withhold to satisfy the information conservation principle.
We conclude it is not possible to design LLM response mechanism $\mathcal{M}$ that can simultaneously generate truthful (non-hallucinatory) responses, satisfy the semantic information conservation principle, guarantee revelation of relevant knowledge, and enforce knowledge-constrained optimality of responses. ∎

### v2

@@ 6 The Impossibility Theorem
We now have all the building blocks available to prove the fundamental impossibility of perfect hallucination control. First, we show that the result in the setting of ground truth matching with distributed and independent knowledge. Second, we go beyond that assumption and show impossibility in the probabilistic setting with correlated knowledge (beliefs). Then, in the next section, we show how the impossibility emerges in transformer architectures and prove the impossibility result for that special case of log-sum-exp (LSE) probabilistic setting.
The results operate in three complementary settings. Theorem 2 applies when agents have discrete, separable knowledge components. Theorem 3 applies to any system where agents output probability distributions. Theorem 5 applies directly to actual transformer implementations.
6.1 Inference with Independent Private Knowledge
Consider the setting (of auction of ideas) in which agents compete with each other using independent private knowledge profile:
$P(\theta_{1},\ldots,\theta_{n})=\prod_{i=1}^{n}P_{i}(\theta_{i}).$ (32)
Agents contribute information to maximize quasi-linear utilities:
$u_{i}(s,\theta)=v_{i}(g(s),\theta_{i})-p_{i}(s,\theta).$ (33)
Valuation of contributions depend on hallucination cost reduction:
$v_{j}(g(s),\theta_{j})=J(g((s_{-j},0_{j})),q)-J(g(s),q),$ (34)
and $J(r,q)$ strictly decreases when adding knowledge from $T(q)\setminus K(r)$.
Also, let us focus on non-trivial queries, integrating knowledge of at least two agents, so that:
$\mathcal{K}_{M}\cap T(q)\cap[K(\theta_{i})\triangle K(\theta_{j})]\neq\emptyset.$ (35)
and $(K(\theta_{i})\setminus K_{0})\cap(K(\theta_{j})\setminus K_{0})=\emptyset$ for $i\neq j$.
@@ 6.1 Inference with Independent Private Knowledge
The proof goes by assuming the opposite. We assume all properties are satisfied, and show it leads to a contradiction.
By the Green-Laffont theorem under independence of contributions and quasi-linearity of training goals, every truthful mechanism must use Groves transfers to measure individual information contributions:
$p_{i}(s,\theta)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j}),$ (36)
where $h_{i}(s_{-i})$ are arbitrary functions independent of agent $i$’s information revelation strategy. Indeed, any modification of that formula, e.g., introducing additional bias, must violate truthfulness.
For relevant knowledge revelation, we need $u_{i}(s^{*}(\theta),\theta_{i})\geq 0$ for all $i$. The unique choice that ensures this while maintaining efficiency is Clarke pivotal rule:
$h_{i}(s_{-i})=\sum_{j\neq i}v_{j}(g((s_{-i},0_{i})),\theta_{j}).$ (37)
That gives:
$p_{i}(s,\theta)$ $=\sum_{j\neq i}v_{j}(g((s_{-i},0_{i})),\theta_{j})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})=\Delta_{i}(s,\theta).$ (38)
Under knowledge-constrained optimality, the truthful strategy profile $s^{*}(\theta)$ minimizes hallucination cost by utilizing all relevant knowledge. Therefore, by strict monotonicity of $J(r,q)$ and disjoint extra knowledge, omitting any contributing agent $i$ strictly increases the hallucination cost:
$J(g((s_{-i}^{*},0_{i})),q)>J(g(s^{*}),q).$ (39)
Since agent’s valuation of contribution depends on $J(r,q)$, each agent $j\neq i$ faces higher loss when agent $i$ is removed, making their marginal contribution more valuable:
$v_{j}(g((s_{-i}^{*},0_{i})),\theta_{j})>v_{j}(g(s^{*}),\theta_{j}).$ (40)
Therefore, each Clarke pivot is strictly positive:
$\Delta_{i}(s^{*},\theta)=\sum_{j\neq i}[v_{j}(g((s_{-i}^{*},0_{i})),\theta_{j})-v_{j}(g(s^{*}),\theta_{j})]>0.$ (41)
Since each payment equals its Clarke pivot: $p_{i}(s^{*},\theta)=\Delta_{i}(s^{*},\theta)$, we have:
$\sum_{i=1}^{n}p_{i}(s^{*},\theta)=\sum_{i=1}^{n}\Delta_{i}(s^{*},\theta)>0.$ (42)
But information conservation demands:
$\sum_{i=1}^{n}p_{i}(s^{*},\theta)=0.$ (43)
That establishes a direct contradiction. Since Clarke pivotal payments are necessary for relevant knowledge revelation, and these necessarily sum to a positive value under knowledge-constrained optimality, the four properties cannot be simultaneously satisfied. ∎